# Exploring the Interior Structure and Mode of Tidal Heating in Enceladus

### v3

@@ 6.1.1 Plausible Interior Parameters and Observables
Using the ranges of $\gamma$, $J_{3}$, MoI, and heat inferred by Park et al. (2024), we also obtain considerably improved constraints on the interior parameters compared to the case where we use broader uncertainties on the observations (See Figures 18 and 18a in Appendix E). However, several model parameters remain weakly constrained by the currently available observations. The ranges of the obtained plausible interior parameters are listed in Table 4. See Appendix E (Figures 18b and 18c) for the PDFs of all the model parameters associated with each of the cases. We further investigate the correlation between total heat and the observables and the interior parameters to which they are sensitive (Figure 9a). The ranges of the effective viscosities of the shell and the core are demonstrated. The total tidal heating is largely determined by the viscosity of the region where the tidal dissipation is dominated. The rigidity of the shell influences the amplitude of both $k_{2}$ and $h_{2}$. We also explore the correlation between the physical properties of the shell (thickness, density, and rigidity) and the observables that are most sensitive to them, i.e., $\gamma$, $J_{3}$, and $k_{2}$ (Figure 9b). Shell thickness and density influence both $\gamma$ and $J_{3}$, although the effect of the density on $J_{3}$ is stronger than on $\gamma$. $D_{shell}$ also strongly correlates with $k_{2}$. The effect of the shell rigidity on the libration amplitude is small compared to the thickness and density which determine the shell’s moment of inertia. This sensitivity suggests that the shell’s response to the libration forces is mainly determined by inertial forces, consistent with Hemingway & Nimmo (2024). The analysis shows that although libration is currently an effective constraint on the shell thickness, for obtaining very high accuracy on the shell and ocean thickness, measurements of the gravity field alongside the shape are much more impactful. Lower sensitivity of libration is largely due to its dependence on the moment of inertia and flexibility of the shell which include effects from thickness, density, and rigidity of the shell.
6.2 Constraining the Interior Structure with Future Observations
[…]
For the case of lower measurement precisions in Table 3, the results show that the suite of observations can constrain the thickness of the shell and the radius of the rocky core to within $\pm$2.7 km, and $\pm$7 km, respectively (3-$\sigma$ confidence level). When tighter constraints on the observables are adopted, the mean thickness of the shell can be constrained within approximately 1.5 km. The ocean density is constrained within $\pm$60 kg/m3 and $\pm$45 kg/m3 in the cases where the tidal dissipation is dominated in the core and the shell, respectively. This low precision is roughly equivalent to a knowledge of seawater salinity to within 40 psu (Millero et al., 2008, Earth’s ocean salinity is 35 psu or parts per thousand). An improved resolution of approximately $\pm$20 kg/m3 can be obtained if the higher precision scenario mentioned in Table 3 is used. The corresponding 20 psu salinity precision would allow for distinguishing the nominal 10 psu salinity of Postberg et al. (2009) from an Earth-like or greater salinity.
The viscosity and rigidity of the core are not well constrained if tidal dissipation is dominated in the shell because the range of $\eta_{\rm core}$ precludes a notable effect on the global tidal response of the core. This finding is consistent with the results of Ermakov et al. (2021) and Genova et al. (2024). The density of the core is constrained within 85 kg/m3 and 75 kg/m3 in case tidal dissipation occurs in the core and in the shell, respectively. If higher-precision measurements are acquired, the density of the core can be constrained within approximately 50 kg/m3.
@@ 6.2 Constraining the Interior Structure with Future Observations
6.3 Effect of Structural Heterogeneity in the Ice Shell
The required measurement precisions computed in Section 6.1 are based on a body with variable properties only in the radial direction (Equations 2.1). Based on shape and gravity observations, Enceladus’s ice shell deviates from spherical symmetry, which in turn can change its tidal response (Běhounková et al., 2017; Berne et al., 2023a, b). Here, we compute the tidal displacements of a heterogeneous ice shell and compare it to the response from a shell with constant thickness with same mean thicknesses. We compute the crustal thickness variations by applying the topography model in Park et al. (2024) (with minor modifications) evaluated up to $l=8$ to the outer surface of geometries and altering the shape of the ice-ocean boundary to mirror the non-hydrostatic components of topography at the outer surface. The relationship between structure at the ice-ocean boundary and surface topography follows a formulation for Airy-type isostasy (Hemingway & Matsuyama, 2017) (see e.g. Equation 19 in Berne et al. (2024)). The shape model suggests variations in the thickness of the shell ranging between 5 and 50 km (Figure 11).
@@ 6.3 Effect of Structural Heterogeneity in the Ice Shell
We assess the effect of the ice shell’s structural heterogeneity on the maximum radial and horizontal tidal deformations on the surface compared to those derived in Section 6.1. The overall maximum surface deformation can differ in comparison to the case of constant thickness shell used in previous studies. We compute the maximum displacements in the radial, northerly, and easterly directions for an ice shell with variable thickness and an ice shell of constant thickness with the same average thicknesses equal to 30 km (Figure 12). In this figure, it is assumed that tidal dissipation occurs in the core and the viscosity of the ice shell is high enough to assume purely elastic behavior. The ice shell with heterogeneous thickness demonstrates variations in the radial and horizontal directions reaching 50 cm compared to a shell of constant thickness (Figure 12). Larger deviations are observed for all components around the South pole, as expected, due to the significant variations in the thickness of the shell (Berne et al., 2023b). The global maximum radial displacement for the case of a shell with variable thickness is approximately 20 cm higher than the constant thickness case. The significant deviation of surface deformation demonstrated in these plots implies that for Enceladus, $h_{2}$ and $l_{2}$ can provide rough estimates of the actual surface displacements as commonly used for spherically symmetric bodies, and for a precise calculation of the displacement at any point on the surface, the effect of structural heterogeneity is important. Since the globally maximum radial and horizontal displacements for a body of constant shell thickness are smaller than those for a shell with thickness variations, the required accuracies presented in Section 6.1 derived by assuming variation only in radial direction are conservative and thus remain valid.
Figure 12: Amplitude of tidal deformation in radial, northerly, and easterly directions for two cases of an ice shell with variable thickness as shown in Figure 11 and a constant shell thickness with the same average thickness of 30 km. Here, it is assumed that the tidal dissipation is happening in the rocky core and the viscosity of ice shell is 1020 Pa.s, implying elastic deformations in the shell. Contours denote intervals of 0.05 m
6.4 Feasibility of the Required Measurements
@@
An onboard altimeter would be ideal for directly measuring radial deformation (e.g., Steinbrügge et al., 2015; Thor et al., 2020). By assuming a similar orbiting geometry and sampling strategy with an altimetry precision of <<5 m, the radial deformation could be recovered with an accuracy <25<25 cm (1-$\sigma$) using surface tie points (i.e. repeat observation of the same surface point at different orbital phases).
@@ 7 Conclusions and discussion
Determining the interior properties of Enceladus remains challenging given the currently available observations. We explored the possibility of determining the internal structure of Enceladus and the tidal heating processes by means of a suite of geodetic measurements. We presented constraints on the interior structure of Enceladus by revisiting the latest analysis of geodetic data obtained by Cassini. Based on these analyses, we provide estimates of the plausible ranges of the currently unavailable tidal response, i.e., $k_{2}$, $h_{2}$, and $l_{2}$, and their associated phase lags. The tidal Love numbers of Enceladus compatible with the currently available observations span a wide range depending on the interior properties. We showed that the potential measurement of the $k_{2}$ and phase lag can be explained by tidal heat generated in the shell or in the rocky core, representing a case of degeneracy that requires further constraints to be resolved. Therefore, neglecting the part of the parameter space associated with core-dominated tidal dissipation can result in an erroneous interpretation of the time-variable gravity field for the interior structure. Our analysis shows that measuring tidal displacements and/or phase lags can discriminate between core-dominated and shell-dominated tidal dissipation cases. If both real and imaginary parts of $k_{2}$ is measured within 0.002, the measurement of the maximum vertical and horizontal displacements within 82 cm and 40 cm are sufficient to obtain a 2-$\sigma$ confidence for discriminating between the two heat generation scenarios. The phase lag associated with $h_{2}$ differs significantly between the two cases of tidal dissipation. In the case of a core-dominated tidal dissipation and an elastic ice shell, $\epsilon_{h_{2}}$ is negligible, while for a shell-dominated tidal dissipation, $\epsilon_{h_{2}}$ can reach tens of degrees (10–50). The requirements for measuring $h_{2}$ and $l_{2}$ are relaxed when improved constraints on $k_{2}$ are obtained. If real and imaginary parts of $k_{2}$ is constrained within 0.0002, measuring the vertical and/or tangential tidal displacements within 116 cm and 51 cm, respectively, is sufficient to discriminate between the two classes of tidal dissipation. We note that while measuring the tidal phase lags is probably necessary to constrain the heat generation mechanisms in the case of Europa and Ganymede as considered by Hussmann et al. (2016), here we demonstrate that in the case of Enceladus, the anelastic deformations can be large enough to constrain the first-order heat generation scenario. Meanwhile, measurements of phase lags for Enceladus can considerably improve our constraints on the rheological properties of the interior.
@@
We also investigated the effect of the structural heterogeneities of the ice shell on the tidal response. Our results show that in the case of a largely elastic shell the surface displacements can significantly deviate from those of a shell of constant thickness. These deviations can reach more than $\sim$50 cm for the thinnest parts of a heterogeneous shell compared to a shell of constant thickness with the same global mean thickness. This observation is important for mission concept design purposes and for determining the locations at which the maximum displacements are measurable. We quantified the effect of a tidally active core on the tidal Love numbers of Enceladus with a heterogeneous ice shell. We found that if the core is tidally active to the extent needed to explain the steady-state heat budget, its effect on the observed $k_{2}$ in the presence of a heterogeneous shell can also be significant.
@@
We conducted inversions using synthetic observables and demonstrated their efficacy for probing the main interior structure parameters such as the rheology, density, and thickness of the layers, using a feasible set of measurements and uncertainties. As one of the important factors in determining habitability, we demonstrated that high- and low-density ocean scenarios differing by more than 40 psu (see Postberg et al., 2009; Glein et al., 2018; Glein & Waite, 2020; Fifer et al., 2022; Postberg et al., 2023) can be distinguished using geodetic measurements. Such different interpretations originate mainly from the underlying assumptions on the provenance of the plume material, the potential change in their composition during their transport from the ocean, and the undetermined extent of overturning circulation and possible ice-ocean interactions (Zhu et al., 2017; Soderlund, 2019; Kang & Flierl, 2020; Lobo et al., 2021; Kang, 2022; Meyer et al., 2025; Ames et al., 2025; Bouffard et al., 2025). This constraint obtained on the density is only marginal, but it is strengthened when tighter constraints from the gravity field are obtained. Constraining core density and rheology can help distinguish between a low-density porous core and a higher-density rocky core and help us understand its composition, formation, and evolution.
@@
We highlight that the measurement requirements presented in this paper are computed under the assumption that the total produced heat produced is partitioned by at least a ratio of 90% and 10% between the core and the shell. These cases are the most likely scenarios, considering the patterns observed in Figure 1 where very limited values of ice and rock viscosities represent comparable dissipation in the shell and in the core. However, in a general case where partitioning of heat between the shell and the core are comparable, quantifying the tidal dissipation rate in the two parts of the body is possible by measuring the complex $k_{2}$ and $h_{2}$, and the density structure of the body (Equation 16). In this case, the uncertainty of quantifying the heat partitioning depends on the uncertainties on the parameters and observables involved in Equation (16). We also note that the computed values of shell and core viscosity should be considered as the “effective” viscosity. The viscosity profiles can depend on the variation of parameters such as porosity, temperature, and the water saturation percentile with depth. However, even if other mechanisms than viscoelastic dissipation are producing the inferred tidal heating, the signature on the tidal response will remain the same. We also point out that the methodology presented here does not address the possibility of tidal dissipation in the ocean, as a result of resonantly enhanced tidal flow as discussed in previous work (Hay & Matsuyama, 2017; Matsuyama et al., 2018; Tyler, 2020). A small $h_{2}$ amplitude and phase lag can rule out the occurrence of tidal dissipation in the shell, but may be associated with tidal dissipation occurring either in the ocean or in the core. However, significant ocean tidal heating at Enceladus is very unlikely to take place because it would require a very thin ocean (less than a few kilometers thick) to cause the resonance (Hay & Matsuyama, 2017; Matsuyama et al., 2018). Such a thin ocean is incompatible with the constraints we obtain on the interior structure from the available estimates of the libration, gravity field, and total mass of Enceladus.
By extending the framework presented in this study to incorporate the future observations of total heat loss and comparison with the total tidally produced heat, we can understand whether Enceladus is in a thermal and orbital steady state, or alternatively, it is experiencing a transient period during which the ocean may become thicker or thinner. The rate of heat loss can be quantified through precise thermal IR or microwave radiometry, or combined constraints on the structure of the shell and its thermal conductivity. The orbital state can be studies by the ground-based or spacecraft astrometry. The suite of needed geodetic measurements can be achieved by a future mission that provides additional constraints on the static gravity field, topography, libration, tidal Love numbers, and rate of surface heat loss.

### v5

@@ 5.3 Constraining the interior of Enceladus with future observations
For the case of lower measurement precisions in Table 3, the results show that the suite of observations can constrain the thickness of the shell and the radius of the rocky core to within $\pm$2.0 km, and $\pm$5 km, respectively (2-$\sigma$ uncertainty). When tighter constraints on the observables are adopted, the mean thickness of the shell can be constrained within approximately 1.2 km. The ocean density is constrained within $\pm$40 kg/m3 and $\pm$30 kg/m3 in the cases where the tidal dissipation is dominated in the core and the shell, respectively. This low precision is roughly equivalent to a knowledge of seawater salinity to within 30 psu (Earth’s ocean salinity is 35 psu or parts per thousand, Millero et al., 2008). An improved resolution of approximately $\pm$15 kg/m3 can be obtained if the higher precision scenario mentioned in Table 3 is used. The corresponding 15 psu salinity precision would allow for distinguishing the nominal 10 psu salinity of Postberg et al. (2009) from an Earth-like or greater salinity.
The viscosity and rigidity of the core are not well constrained if tidal dissipation is dominated in the shell because the range of $\eta_{\rm core}$ precludes a notable effect on the global tidal response of the core. This finding is consistent with the results of Ermakov et al. (2021) and Genova et al. (2024). The density of the core is constrained within 65 kg/m3 and 50 kg/m3 in case tidal dissipation occurs in the core and in the shell, respectively. If higher-precision measurements are acquired, the density of the core can be constrained within approximately 35 kg/m3.
@@
5.4 Structural heterogeneity in the ice shell
The required measurement precisions computed in Section 5.2 are based on a body without any lateral variation of properties (equations 2.1). However, shape and gravity observations suggest that Enceladus’s ice shell deviates from spherical symmetry, which in turn can change its tidal response (Běhounková et al., 2017; Berne et al., 2023a, b). Here, we compute the tidal displacements of a heterogeneous ice shell and compare it to the response of a shell with a constant thickness. The two cases have the same mean thicknesses. We compute the crustal thickness variations by applying a modified topography model from Park et al. (2024) evaluated up to $l=8$ to the outer surface of geometries and altering the shape of the ice-ocean boundary to mirror the non-hydrostatic components of topography at the outer surface. The topography model presented in Park et al. (2024) contained artifacts including a depression and a mound, as a result of ignoring to incorporate the available limb data. The present analysis is based on this updated shape model. An erratum by Park et al. (2024) is in preparation to publish the revised shape model. The relationship between structure at the ice-ocean boundary and surface topography follows an Airy-type isostasy (Hemingway & Matsuyama, 2017) (see e.g. equation 19 in Berne et al. (2024)). For this shape model, the thickness of the shell ranges between 5 and 50 km (Figure 11).
@@ 5.4 Structural heterogeneity in the ice shell
We assess the effect of the ice shell’s structural heterogeneity on the maximum radial and horizontal tidal deformations on the surface compared to those derived in Section 5.2. The overall maximum surface deformation can differ in comparison to the case of constant thickness shell used in previous studies. We compute the maximum displacements in the radial, northerly, and easterly directions for an ice shell with variable thickness and an ice shell of constant thickness with the same average thicknesses equal to 30 km (Figure 12). In this figure, we assume that tidal dissipation occurs in the core and the viscosity of the ice shell is high enough to assume a purely elastic behavior. The ice shell with a heterogeneous thickness demonstrates variations in the radial and horizontal directions reaching 50 cm compared to a shell of constant thickness (Figure 12). Larger deviations are observed for all components around the South pole, as expected, due to the significant thickness variations from the mean thickness (Berne et al., 2023b). The global maximum radial displacement for the case of a shell with variable thickness is approximately 20 cm higher than the constant thickness case. The significant deviation of surface deformation demonstrated in these plots implies that for Enceladus, $h_{2}$ and $l_{2}$ can provide rough estimates of the actual surface displacements as commonly used for spherically symmetric bodies, and for a precise calculation of the displacement at any point on the surface, the effect of structural heterogeneity has to be accounted for. Since the globally maximum radial and horizontal displacements for a body of constant shell thickness are smaller than those for a shell with thickness variations, the required accuracies presented in Section 5.2 derived by assuming variation only in radial direction are conservative and thus remain valid.
Figure 12: Amplitude of tidal deformation in radial, northerly, and easterly directions for two cases of an ice shell with variable thickness as shown in Figure 11 and a constant shell thickness with the same average thickness equal to 30 km. Here, it is assumed that the tidal dissipation is happening in the rocky core and the viscosity of ice shell is 1020 Pa.s, implying elastic deformations in the shell. Contours denote intervals of 0.05 m
5.5 Feasibility of required measurements
@@
An onboard altimeter would be ideal for directly measuring radial deformation (e.g., Steinbrügge et al., 2015; Dirkx et al., 2019; Thor et al., 2020). By assuming a similar orbiting geometry and sampling strategy with an altimetry precision of <<5 m, the radial deformation could be recovered with an accuracy <25<25 cm (1-$\sigma$) using surface tie points (i.e. repeat observation of the same surface point at different orbital phases).
@@ 6 Conclusions and discussion
Determining the interior properties of Enceladus remains challenging given the currently available observations. We evaluated the extent to which geodetic measurements can constrain Enceladus’s interior structure and the mode of its tidal heating. We presented constraints on the interior structure of Enceladus by revisiting the latest analysis of geodetic data acquired by Cassini. Based on these analyses, we provide estimates of the plausible ranges of the currently unavailable tidal response, i.e., $k_{2}$, $h_{2}$, and $l_{2}$, and their phase lags. The tidal Love numbers of Enceladus compatible with the currently available observations span a wide range depending on the interior properties. We showed that a potential measurement of the $k_{2}$ and phase lag can be explained by tidal heat generated either in the shell or in the core, representing a case of degeneracy that requires additional constraints to be resolved. Our analysis shows that measuring either the radial or tangential tidal displacements can discriminate between core-dominated and shell-dominated tidal dissipation cases. Measuring the associated phase lags can also differentiate between the two scenarios. If both real and imaginary parts of $k_{2}$ is measured within 0.002, measuring the maximum vertical and horizontal displacements within 82 cm and 40 cm are sufficient to obtain a 2-$\sigma$ confidence for discriminating between the two heat generation scenarios. The phase lags of $h_{2}$ and $l_{2}$ differ significantly between the two cases of tidal dissipation. In the case of a core-dominated tidal dissipation and an elastic ice shell, both $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ are very small, while for a shell-dominated tidal dissipation, both $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ can reach tens of degrees (10–50). The requirements for measuring $h_{2}$ and $l_{2}$ are further relaxed in case refined constraints on $k_{2}$ are obtained. If real and imaginary parts of $k_{2}$ is constrained within 0.0002, measuring the vertical and/or tangential tidal displacements within 116 cm and 51 cm, respectively, is sufficient to discriminate between the two classes of tidal dissipation. We note that while in the case of most of the moons such as Europa and Ganymede measuring the tidal phase lags is probably necessary to constrain the heat generation mechanisms (Hussmann et al., 2016), here we demonstrate that the anelastic deformations at Enceladus can be large enough to differentiate between the main heat generation scenarios. Yet, obtaining tighter constraints on the rheology of the interior layers demands measuring the phase lags, in addition to the amplitudes of the Love numbers.
@@
We also investigated the effect of the structural heterogeneities of the ice shell on the tidal response. Our results show that in the case of a largely elastic shell the surface displacements can significantly deviate from those of a shell of constant thickness. These deviations can reach more than $\sim$50 cm for the thinnest parts of a heterogeneous shell compared to a shell of constant thickness with the same global mean thickness. This observation is important for mission concept design purposes and for determining the locations at which the displacements are measurable. We quantified the effect of a tidally active core on the tidal Love numbers of Enceladus with a heterogeneous ice shell. We found that if the core is tidally active to the extent needed to explain the steady-state heat loss rate, its effect on the observed $k_{2}$ in the presence of a heterogeneous shell can be comparable or larger than the effect of the ice shell heterogeneities.
@@
We conducted inversions using synthetic observables and demonstrated their efficacy for probing the main interior structure parameters, using a set of feasible measurements and uncertainties. As one of the important factors in determining habitability, we demonstrated that high- and low-density ocean scenarios can be differentiated using the geodetic measurements. Such different interpretations originate mainly from the underlying assumptions on the provenance of the plume material, the potential change in their composition during their transport from the ocean, and the undetermined extent of overturning circulation and possible ice-ocean interactions (Zhu et al., 2017; Soderlund, 2019; Kang & Flierl, 2020; Lobo et al., 2021; Kang, 2022; Meyer et al., 2025; Ames et al., 2025; Bouffard et al., 2025). These interpretations differ by more than 40 psu (see Postberg et al., 2009; Glein et al., 2018; Glein & Waite, 2020; Fifer et al., 2022; Postberg et al., 2023). The constraint obtained on the ocean density from our geophysical measurements is not strong, but it can differentiate between a low-density nearly pure ocean and a high-salinity high-density ocean. This constraint can be strengthened when precise measurements, mainly from the static and time variable gravity field, are obtained. We also constrain the core density and rheology which can help distinguish between a low-density porous core and a higher-density rocky core and help us understand its composition, formation, and thermo-chemical evolution.
The analyses presented here imply underlying assumptions as elaborated here. We base the estimated tidal dissipation used in our analysis on the heat loss rates from previous studies (i.e., Hemingway & Mittal (2019); Howett et al. (2011); Nimmo et al. (2018); Howett et al. (2025)). Assuming that the estimated heat loss is equal to the internal heat production by tides implies that Enceladus is currently in thermal steady state. However, the thermal state and the orbital state of Enceladus, which tightly affect one another, are not well understood. If Enceladus is currently experiencing an episodic or periodic cycle of melting and freezing of its hydrosphere, as suggested by O’Neill & Nimmo (2010); Nimmo et al. (2018); Goldreich et al. (2025), the rates of heat production and the heat loss can be different. Although, these scenarios are hindered by requiring that we happen to be observing Enceladus in a special time of its overall evolution. Alternatively, the resonance locking hypothesis implies a nearly steady state and that the high tidal dissipation at Enceladus is sustained over long timescales. This scenario can best explain the observed Saturnian system as discussed in detail in Nimmo et al. (2018, 2023). Through obtaining independent observations of total heat loss and tidally produced heat and comparing them, we can understand whether Enceladus is in a thermal and orbital steady state.
@@
We also note that he measurement requirements presented in this paper are computed under the assumption that the total heat produced is partitioned by at least a ratio of 90% and 10% between the core and the shell. These cases are the most likely scenarios, considering the patterns observed in Figure 2 where very limited values of ice and rock viscosities represent comparable dissipation in the shell and in the core. However, in a general case where partitioning of heat between the shell and the core are comparable, quantifying the tidal dissipation rate in the two parts of the body is possible by measuring the complex $k_{2}$ and $h_{2}$, and the density structure of the body (equation 11 derived here). In this case, the uncertainty of quantifying the heat partitioning depends on the uncertainties on the parameters and observables involved in equation (11).
We also point out that the computed values of shell and core viscosity should be considered as the “effective” viscosity of the layer. The viscosity profiles can depend on the variation of parameters such as porosity, temperature, and the water saturation percentile with depth. However, even if other mechanisms than viscoelastic dissipation are producing the inferred tidal heating, their effect on the tidal response will remain the same, and thus our analysis remains valid.
The methodology presented here does not consider the possibility of tidal dissipation in the ocean, as a result of resonantly enhanced tidal flow as discussed in Hay & Matsuyama (2017); Matsuyama et al. (2018); Tyler (2020). A small $h_{2}$ amplitude and phase lag can rule out the occurrence of tidal dissipation in the shell, but may be associated with tidal dissipation occurring either in the ocean or in the core. However, significant ocean tidal heating at Enceladus is very unlikely because it would require a very thin ocean (less than a few kilometers thick) to cause the resonance (Hay & Matsuyama, 2017; Matsuyama et al., 2018). Such a thin ocean is incompatible with the constraints we obtain on the interior structure from the available estimates of the libration, gravity field, and total mass of Enceladus.
The necessary measurements suggested here can be achieved by a future mission that provides additional constraints on the static gravity field, topography, libration, tidal Love numbers, and rate of surface heat loss. The orbital state can be studied by the ground-based or spacecraft astrometry. The methodology presented here is general and can be used to investigate the modes of tidal heating in other icy satellites such as Europa, Ganymede, and Titan, through geophysical data which will become available by the upcoming Europa Clipper, Juice, and Dragonfly missions.